OpenStudy (anonymous):

Which line is parallel to the line that passes through the points (-1, 2) and (5, -4)? y= -x+1 y= x+3 y= 1/3x+4 y= -1/3x-4

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@pooja195

OpenStudy (owlcoffee):

Since you are given the two points belonging to the line that is parallel to the solution one, I would suggest you first find the equation of the line composed by those two points and firstly calculating the slope composed by it. Let's call the line composed by A(-1,2) and B(5,-4) as line "r", we first find the slope of this line using the formula: \[m=\frac{ y_B -y_A}{ x_B-x_A }\] Thereby: \[m_r=\frac{ -4-2 }{ 5-(-1) }\] \[m_r = \frac{ -6 }{ 6 } \rightarrow m_r=-1\] Now, there let's call the line that is parallel to "r" as "t" and mention the sufficient condition for the parallelism of two lines, which is that their slopes must be equal: \(m_r = m_t\). So, following the structure of the equation of the line \[y=m_t x +c\]. Which line of the given solutions does have a slope of "-1"?

OpenStudy (anonymous):

y= -x+1??

OpenStudy (owlcoffee):

correct.